OpenStudy (anonymous):

Multiply: 1.) [sqrt.2x^2] * [10x^6] 2.) [2x sqrt.12x] * [3 sqrt.4x] 3.) 6 sqrt.5 ( [srt.11] + [sqrt.5] )

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

It's impossible to solve it. You need to make it more clear. Try using the equation button or taking a picture of the questions and attaching it here.

OpenStudy (anonymous):

^ = to the power sqrt. = that square root sign

OpenStudy (anonymous):

oh nevermind i just saw the eqaution button , holdon

OpenStudy (anonymous):

Yeah, but would it be \[\sqrt{(2x)^{2}} * (10x)^{^{6}}\] or \[(\sqrt{2x})^{^{2}}*10x ^{6}\] ?

OpenStudy (anonymous):

\[1.) \sqrt{2x ^{2}} \times \sqrt{10x ^{6}}\]

OpenStudy (anonymous):

2.) \[(2x \sqrt{12x} ) (3\sqrt{4x})\]

OpenStudy (anonymous):

3.) \[6\sqrt{5} (5\sqrt{11} + \sqrt{5})\]

OpenStudy (anonymous):

\[\sqrt{2} x * \sqrt{10} x ^{3}\] \[2 \sqrt{5} x ^{4}\]

OpenStudy (anonymous):

ok now?

OpenStudy (anonymous):

why wouldn't it be \[2x ^{2}\sqrt{5}\]

OpenStudy (anonymous):

nevermind, I thought the x^4 was in the sqrt.

OpenStudy (anonymous):

well khmath i think your right

OpenStudy (anonymous):

oh never mind neither answers show here

OpenStudy (anonymous):

would it be \[2x ^{4} \sqrt{5}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

and for the rest of the problems ?

OpenStudy (anonymous):

u figured out the second one ...right??/

OpenStudy (anonymous):

it wasnt right

OpenStudy (anonymous):

i just was asking if it was

OpenStudy (anonymous):

ok lets do it again...

OpenStudy (anonymous):

are you sure you know how to do this ?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

ok just making sure

OpenStudy (anonymous):

4sqrt(3) x * sqrt(x) * 6 * sqrt(x) 24 sqrt(3) *x

OpenStudy (anonymous):

mm.. 24 sqrt(3) *x * sqrt(x)

OpenStudy (anonymous):

for the second ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

no it doesnt give that answer here as an option :/

OpenStudy (anonymous):

wait i dont get which one you wrote

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

but which was your answer

OpenStudy (anonymous):

let me rewrite again

OpenStudy (anonymous):

one min

OpenStudy (anonymous):

24 sqrt3 x^2

OpenStudy (anonymous):

final answer...

OpenStudy (anonymous):

now first two are solved...okay?

OpenStudy (anonymous):

\[24\sqrt{3x ^{2}}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no ...x^2 is not under the square root

OpenStudy (anonymous):

so just 3x alone ?

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

\[24 \sqrt{3} \]\[x ^{2}\]

OpenStudy (anonymous):

they r together

OpenStudy (anonymous):

x goes before where ?

OpenStudy (anonymous):

doesnt matter where u write x ,

OpenStudy (anonymous):

thanks , but nevermind i really dont think you know how to this . to many attempts , but thanks for trying to help